OpenStudy (anonymous):

integer exponents 4+(-6)^0 -4^-1

OpenStudy (anonymous):



OpenStudy (anonymous):

please help :(

OpenStudy (calculusxy):

Do you wish to evaluate this?

OpenStudy (anonymous):

what evaluate?

OpenStudy (calculusxy):

Meaning like simplify

OpenStudy (anonymous):

anything to the zero power = 1. a negative power is just the absolute value but in the numerator. This gives you: 4 + 1 - 1/4

OpenStudy (anonymous):

oh ya :D

OpenStudy (anonymous):

oops line 2 i meant denominator

OpenStudy (anonymous):

so that means anything that has 1 over it is 1 then that number in fraction? or evertyhting even if it had say 5 ?

OpenStudy (calculusxy):

Easier to say, what does 1/4 equal to when it is in decimal?

OpenStudy (anonymous):

if you have a negative exponent, you make it positive and put number being raised in denominator. \[2^{-5} = \frac{ 1 }{ 2^{5} }\]

OpenStudy (anonymous):

0.25?

OpenStudy (anonymous):

oh but i need to make it decimal

OpenStudy (anonymous):

plug into calculator

OpenStudy (calculusxy):

Yes. Since it is negative 1/4, it will be -0.25. The expression will then be 4+(-6)-(-0.25) Then you just simplify it.

OpenStudy (anonymous):

u need to simplify it even more?

OpenStudy (calculusxy):

No, you have to simplify the given expression (that I provided).

OpenStudy (anonymous):

isnt the answer -0.25?

OpenStudy (calculusxy):

No. First, you do the addition, what do you then get?

OpenStudy (anonymous):

we never did that we only get the decimal and thats all

OpenStudy (calculusxy):

I hope that you know that we are using the order of operations.

OpenStudy (calculusxy):

Are you aware of the order of operations?

OpenStudy (anonymous):

no -0.25 is just the answer to -4^-1. You are trying to solve 4+(-6)^0 -4^-1 so you need 4 + 1 and then -0.25

OpenStudy (anonymous):

we arentt supposed to do that ur confusing Me >.<

OpenStudy (anonymous):

4+1=5

OpenStudy (calculusxy):

P for parenthesis E for exponents M for multiplication D for division A for addition A for subtraction

OpenStudy (calculusxy):

I meant S

OpenStudy (anonymous):

5 - 0.25 = 4.75

OpenStudy (calculusxy):

4-(-6)=-2 -2+(0.25)=-2.25

OpenStudy (anonymous):

OHH ok

OpenStudy (anonymous):

@piglet9

OpenStudy (anonymous):

@calculusxy I think you got that wrong there buddy... (-6)^0 is not -6

OpenStudy (anonymous):

how about this.

OpenStudy (anonymous):

\[3(-9)^{0}+4^{-2} \]

OpenStudy (anonymous):

3(1) + 1/16 = 3.0625

OpenStudy (anonymous):

isnt it 0.0625?

OpenStudy (anonymous):

1/16 is 0.0625 but you have to add the 3

OpenStudy (anonymous):

oh ya in the begging one right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how about this

OpenStudy (anonymous):

\[6+(-5)^{-2} -(4=3)^{0} \]

OpenStudy (anonymous):

I don't think you can have parentheses across the equals sign like that...

OpenStudy (anonymous):

my bad its supposed to be +